OpenStudy (anonymous):

2. The height of one solid limestone square pyramid is 24 m. A similar solid limestone square pyramid has a height of 30 m. The volume of the larger pyramid is 16,000 m3. Determine each of the following, showing all your work and reasoning. a) The scale factor of the smaller pyramid to the larger pyramid in simplest form. b) The ratio of the area of the base of the smaller pyramid to the larger pyramid. c) Ratio of the volume of the smaller pyramid to the larger. d) The volume of the smaller pyramid.

OpenStudy (anonymous):

Please any help will be greatly appreciated

OpenStudy (anonymous):

i assume the base areas are equal as no info otherwise the only difference is the height...so scale factor is 24/30=4/5 vol.1/3 base area(height)so as base area the same vol in ratio 4/5 smaller vol.=4/5of larger =4/5*16000=12800